Jack's overtime wage is $3 per hour more than his regular hourly wage. He worked for 5 hours at his regular wage and 4 hours at the overtime wage. He earned $66.00 Find his regular wage.
Answer by checkley75(3666)

5x+4(x+3)=66
5x+4x+12=66
9x=66-12
9x=54
x=54/9
x=$6 for regular hours thus the overtime hourly rate=6+3=$9 for the overtime hours.
5*6+4*9=66
30+36=66
66=66